# Open-Source Highly Parallel Electromagnetic Simulations for Superconducting Circuits

### v1

@@ IV.2 Qubits
After validating SQDMetal’s accuracy for resonator frequencies, we now compare its predictions for transmon qubit parameters against experimental measurements. In this comparison, we examine nine transmon qubits from two separate devices. The first device, labeled Device A, fabricated by Gaikwad et al. [27], contains three transmon qubits. The second device, Device B, consists of six transmon qubits fabricated by MIT Lincoln Laboratory and was used in the work by Shanto et al. [15]. For each qubit in both devices, we compared the simulated and measured values for the qubit transition frequency $f_{ge}$, qubit anharmonicity $\alpha_{q}/2\pi$, readout resonator frequency $f_{r}$, and coupling strength between the qubit and readout resonator $g/2\pi$. These results are summarized in Table 4.
Table 4: Comparison of simulated and experimental transmon parameters. This table lists the qubit frequency ($f_{ge}$), qubit anharmonicity ($\alpha_{q}$), resonator frequency ($f_{r}$), and coupling strength ($g$) obtained from simulation and experiment for each device. Simulated values are shown to the left of the measured values, with measured data given in square brackets. Qubits from Device A [27] and Device B [15] are labeled with the prefixes A and B, respectively.
@@
The energy participation ratio (EPR) method was used to compute the qubit parameters [10]. The EPR approach offers several advantages over other methods, namely, it requires only a single eigenmode simulation and it achieves good accuracy by using Maxwell’s curl equations to perform full-wave field calculations. Full-wave simulations are advantageous as they capture high-frequency effects, such as equivalent series inductance (ESL), which is often neglected in simplified models.
For Device B, the Josephson energy $E_{J}$ was extracted from experimental measurements of the qubit frequency and anharmonicity using the expression [25]:
@@
where $\omega_{q}=2\pi f_{ge}$ is the qubit angular frequency and $E_{C}$ is the charging energy, which is approximately equal to qubit anharmonicity in the transmon regime ($E_{J}\gg E_{C}$) [28].
For Device A, applying Eqn. (2) directly to the measured values of $\omega_{q}$ and $\alpha_{q}$ produced $E_{J}$ values which were overestimated, leading to inaccurate simulation results. To improve agreement, $E_{J}$ was recalculated using Eqn. (2) and the simulated values reported in [27]. This adjustment illustrates the EPR method’s sensitivity to the choice of $E_{J}$ and highlights the importance of accurate experimental extraction of $E_{J}$ to produce consistent simulation results [10, 29]. Once an estimate for $E_{J}$ is obtained, then the corresponding Josephson inductance is calculated from:
@@
where $\phi_{0}=\hbar/2e$ is the reduced flux quantum. For weakly nonlinear circuits operated in the dispersive regime, $|\omega_{r}-\omega_{q}|=\Delta\gg g$, such as a transmon capacitively coupled to a resonator, the Hamiltonian parameters can be derived directly from the EPRs [10]. The anharmonicity of the qubit and resonator, along with the dispersive shift between them, can be expressed in terms of the participation ratios as follows:
@@
where $p_{q}$ and $p_{r}$ are the participation ratios of the qubit and resonator modes, respectively, and $\omega_{q}$ and $\omega_{r}$ are their corresponding angular frequencies. These expressions are used in this study to compute the qubit parameters in Table 4. However, these expressions are only appropriate for qubits operating in both the dispersive and transmon regimes. For qubits operating outside these regimes numerical diagonalization, as outlined in Appendix B, should be used. For reference, an explicit example of the calculation of transmon parameters using Eqns. (4), (5) and (6) is provided in Appendix C.
Once the dispersive shift, $\chi_{qr}$, is determined, the coupling strength $g$ can be obtained from the following expression [15]:
@@
where $\Delta=\omega_{r}-\omega_{q}$ and $\Sigma=\omega_{r}+\omega_{q}$. Eqn. (7) is proposed to be more accurate for calculating $\chi_{qr}$ because it is not derived using the rotating wave approximation (RWA), which can lead to inaccuracies due to ignoring the faster rotating terms [15].
To evaluate the accuracy of the simulated results, we calculated the root-mean-square error (RMSE) for each qubit parameter, both in absolute terms and as a percentage, using the data presented in Table 4. The resulting values, summarized in Table 5, show reasonable agreement between simulation and experiment. The percentage RMSE values for the qubit and resonator frequencies are 5.99% and 3.54%, respectively, while the anharmonicity and coupling strength exhibit larger deviations of 13.25% and 24.53%. These discrepancies likely stem from simplified material and geometric assumptions in the simulations, as well as from the assumption that device fabrication was performed with minimal imperfections. In particular, the coupling strength $g$ is highly sensitive to fabrication variations such as etch bias [5].
The differences observed for $\alpha_{q}$ and $g$ in this study are consistent with values reported in other studies using the EPR method. For example, Gaikwad et al. [27], who fabricated Device A, presented simulated qubit parameters from which percentage RMSE values of 17.21% and 16.40% were determined for $\alpha_{q}$ and $g$, respectively. Similarly, Yuan et al. [29] found a 13.5% difference for $\alpha_{q}$ when using the EPR method. Improved agreement between simulation and experiment is expected by incorporating additional physical effects, such as kinetic inductance [30], and by modeling the full 3D device geometry, which more accurately captures fringing fields and current distributions.
@@ VI Acknowledgments
We would like to acknowledge useful discussions with Eli Levison-Falk, Saikat Das and Sadman Ahmed Shanto from the Levinson-Falk Lab (LFL) at the University of Southern California. Moreover, we would also like to thank the LFL for providing measurement data and Qiskit Metal designs for superconducting qubits, allowing the completion of the experimental results section of the paper. In addition, we would also like to thank Ahbishek Chakraborty from Chapman University who organised the collaboration with LFL and provided useful feedback on SQDMetal. We would like to acknowledge and thank Marlies Hankel and David Green from the research computing centre (RCC) at the University of Queensland (UQ). They provided valuable assistance, helping to install and run Palace on the Bunya cluster at UQ. Finally, we would like to thank Hugh Carson and Andrew Keller from AWS Palace who provided useful feedback on the mesh convergence study and provided helpful insight into Palace.
@@ Appendix B Energy Participation Ratio Method
A superconducting quantum circuit can be described as a largely linear electromagnetic environment containing one or more embedded nonlinear elements. In most cases, these nonlinear elements are Josephson junctions, although other components such as superconducting nanowires (nanoinductors) [32] may also be used. The core idea of the energy participation ratio (EPR) method is to separate each nonlinear element into its linear and nonlinear constituents, a process known as linearization, and then remove the nonlinear part from the remaining linear circuit. An eigenmode simulation of this linearized structure is then performed to obtain the electromagnetic field distributions. From these results, the EPR is computed, quantifying the fraction of the total electromagnetic energy stored in each nonlinear element. This quantity links the linear eigenmodes of the circuit to its nonlinear dynamics and provides a direct route to determine the contribution of each nonlinear element to the system Hamiltonian. The participation ratio, which serves as the bridge between linear and nonlinear circuit components, will be discussed in greater detail in the following sections.
@@ B.2 Full Hamiltonian of the System
where $m$ represents the electromagnetic mode, $\omega_{m}$ are the resonant frequencies of the modes and $\hat{a}_{m}^{{\dagger}}$ and $\hat{a}_{m}$ are the creation and annihilation operators for the given mode, respectively. If we are using linear elements in our circuit we can always describe the total energy with the Hamiltonian in Eqn. (13). Essentially, for every radiation field mode in the circuit we are associating with it a harmonic oscillator. Note that we have dropped the zero-point energy term, $\hbar\omega_{m}/2$, from Eqn (13) as is customary in renormalization to avoid infinite energies. To express the full Hamiltonian we combine the linear and nonlinear parts as follows:
@@
where the first term, introduced in Eqn. (13), includes the linearized Josephson junction. The second term represents the nonlinear contributions from the total number of Josephson junctions, $J$, present in the circuit. The nonlinear part of the full Hamiltonian can be expressed as:
$\mathcal{E}_{j}(\hat{\phi}_{j})^{\text{nlin}}$ $=-\frac{E_{J}}{24}\hat{\phi_{j}}^{4}+\frac{E_{J}}{720}\hat{\phi}_{j}^{6}+$... (15)
@@
where the angular brackets signify indicate to take the expectation value of the expression. Note that due to the energy balance between the inductive energy and capacitive energy stored in a given mode, the total inductive energy of the mode can be expressed as half the total energy. Then performing the calculation, where we take the expectation value with reference to a given Fock state $|n\rangle$, we end up with the result:
@@
where $S_{mj}\in\{-1,1\}$ is the sign of the junction and corresponds to the direction of current flow across the junction. From Eqn. (21) it can be seen that the participation ratio links the quantum properties of the circuit to its classical electromagnetic behaviour. Furthermore, all the components of the Hamiltonian are now fully specified and the system can now be solved by numerical diagonalization. What has been described above is a brief summary of the work in Ref. [10], for further details please see this paper.
@@ Appendix C Example: Transmon Coupled to a Resonator
The energy participation ratios for each mode are automatically calculated in post-processing by Palace. Recall that the EPR for a given mode is defined as the inductive energy stored in the junction divided by the total inductive energy in the mode. Additionally, the energy balance for a resonantly excited mode, $m$, due to time-averaging is equally split between inductive energy and electric energy [10]:

### v2

@@ IV.2 Qubits
After validating SQDMetal’s accuracy for resonator frequencies, we now compare its predictions for transmon qubit parameters against experimental measurements. In this comparison, we examine nine transmon qubits from two separate devices. The first device, labeled Device A, fabricated by gaikwad_entanglement_2024, contains three transmon qubits. The second device, Device B, consists of six transmon qubits fabricated by MIT Lincoln Laboratory and was used in the work by shanto_squadds_2024. For each qubit in both devices, we compared the simulated and measured values for the qubit transition frequency $f_{ge}$, qubit anharmonicity $\alpha_{q}/2\pi$, readout resonator frequency $f_{r}$, and coupling strength between the qubit and readout resonator $g/2\pi$. These results are summarized in Table 4.
Table 4: Comparison of simulated and experimental transmon parameters. This table lists the qubit frequency ($f_{ge}$), qubit anharmonicity ($\alpha_{q}$), resonator frequency ($f_{r}$), and coupling strength ($g$) obtained from simulation and experiment for each device. Simulated values are shown to the left of the measured values, with measured data given in square brackets. Qubits from Device A [gaikwad_entanglement_2024] and Device B [shanto_squadds_2024] are labeled with the prefixes A and B, respectively.
@@
We used the energy participation ratio (EPR) method was to compute the qubit parameters [minev_energy-participation_2021]. The EPR approach offers several advantages over other methods, namely, it requires only a single eigenmode simulation and it achieves good accuracy by using Maxwell’s curl equations to perform full-wave field calculations. Full-wave simulations are advantageous as they capture high-frequency effects, such as equivalent series inductance (ESL), which is often neglected in simplified models.
For Device B, the Josephson energy $E_{J}$ was extracted from experimental measurements of the qubit frequency and anharmonicity via [blais_circuit_2021]:
@@
where $\omega_{q}=2\pi f_{ge}$ is the qubit angular frequency and $E_{C}$ is the charging energy, which is approximately equal to qubit anharmonicity in the transmon regime ($E_{J}\gg E_{C}$) [koch_charge-insensitive_2007].
For Device A, applying Eqn. 2 directly to the measured values of $\omega_{q}$ and $\alpha_{q}$ produced $E_{J}$ values which were overestimated, leading to inaccurate simulation results. To improve agreement, $E_{J}$ was recalculated using Eqn. 2 and the simulated values reported in [gaikwad_entanglement_2024]. This adjustment illustrates the EPR method’s sensitivity to the choice of $E_{J}$ and highlights the importance of accurate experimental extraction of $E_{J}$ to produce consistent simulation results [minev_energy-participation_2021, yuan_comparison_2022]. Once an estimate for $E_{J}$ is obtained, then the corresponding Josephson inductance is calculated from:
@@
where $\phi_{0}=\hbar/2e$ is the reduced flux quantum. For weakly nonlinear circuits operated in the dispersive regime, $|\omega_{r}-\omega_{q}|=\Delta\gg g$, such as a transmon capacitively coupled to a resonator, the Hamiltonian parameters can be derived directly from the EPRs [minev_energy-participation_2021]. The anharmonicity of the qubit and resonator, along with the dispersive shift between them, can be expressed in terms of the participation ratios as follows:
@@
where $p_{q}$ and $p_{r}$ are the participation ratios of the qubit and resonator modes, respectively, and $\omega_{q}$ and $\omega_{r}$ are their corresponding angular frequencies. These expressions are used in this study to compute the qubit parameters in Table 4. However, these expressions are only appropriate for qubits operating in both the dispersive and transmon regimes. For qubits operating outside these regimes numerical diagonalization, as outlined in Appendix B, should be used. For reference, an explicit example of the calculation of transmon parameters using Eqns. 4, 5 and 6 is provided in Appendix C.
Once the dispersive shift, $\chi_{qr}$, is determined, the coupling strength $g$ can be obtained from the following expression [shanto_squadds_2024]:
@@
where $\Delta=\omega_{r}-\omega_{q}$ and $\Sigma=\omega_{r}+\omega_{q}$. Eqn. 7 is proposed to be more accurate for calculating $\chi_{qr}$ because it is not derived using the rotating wave approximation (RWA), which can lead to inaccuracies due to ignoring the faster rotating terms [shanto_squadds_2024].
To evaluate the accuracy of the simulated results, we calculated the root-mean-square error (RMSE) for each qubit parameter, both in absolute terms and as a percentage, using the data presented in Table 4. The resulting values, summarized in Table 5, show reasonable agreement between simulation and experiment. The percentage RMSE values for the qubit and resonator frequencies are 5.99% and 3.54%, respectively, while the anharmonicity and coupling strength exhibit larger deviations of 13.25% and 24.53%. These discrepancies likely stem from simplified material and geometric assumptions in the simulations, as well as from the assumption that device fabrication was performed with minimal imperfections. In particular, the coupling strength $g$ is highly sensitive to fabrication variations such as etch bias [levenson-falk_review_2024].
The differences observed for $\alpha_{q}$ and $g$ in this study are consistent with values reported in other studies using the EPR method. For example, gaikwad_entanglement_2024, who fabricated Device A, presented simulated qubit parameters from which percentage RMSE values of 17.21% and 16.40% were determined for $\alpha_{q}$ and $g$, respectively. Similarly, yuan_comparison_2022 found a 13.5% difference for $\alpha_{q}$ when using the EPR method. Improved agreement between simulation and experiment is expected by incorporating additional physical effects, such as kinetic inductance [park_improving_2024], and by modeling the full 3D device geometry, which more accurately captures fringing fields and current distributions.
@@ VI Acknowledgments
We would like to acknowledge useful discussions with Eli Levenson-Falk, Saikat Das and Sadman Ahmed Shanto from the Levenson-Falk Lab (LFL) at the University of Southern California. Moreover, we would also like to thank the LFL for providing measurement data and Qiskit Metal designs for superconducting qubits, allowing the completion of the experimental results section of the paper. In addition, we would also like to thank Ahbishek Chakraborty from Chapman University who organised the collaboration with LFL and provided useful feedback on SQDMetal. We would like to acknowledge and thank Marlies Hankel and David Green from the research computing centre (RCC) at the University of Queensland (UQ). They provided valuable assistance, helping to install and run Palace on the Bunya cluster at UQ. Finally, we would like to thank Hugh Carson and Andrew Keller from AWS Palace who provided useful feedback on the mesh convergence study and provided helpful insight into Palace.
Figure 5: Electric field visualization of the fundamental modes for each design. ParaView is used to visualize the results of the eigenmode simulation. (a) Single CPW resonator. (b) Single CPW resonator capacitively coupled to a feedline. (c) Transmon capacitively coupled to a CPW resonator with the top and bottom figures showing the transmon and resonator qubit modes. (d) Two capacitively coupled transmons with the top and bottom figures showing the left and right transmon modes.
@@ Appendix B Energy Participation Ratio Method
A superconducting quantum circuit can be described as a largely linear electromagnetic environment containing one or more embedded nonlinear elements. In most cases, these nonlinear elements are Josephson junctions, although other components such as superconducting nanowires (nanoinductors) [annunziata_tunable_2010] may also be used. The core idea of the energy participation ratio (EPR) method is to separate each nonlinear element into its linear and nonlinear constituents, a process known as linearization, and then remove the nonlinear part from the remaining linear circuit. An eigenmode simulation of this linearized structure is then performed to obtain the electromagnetic field distributions. From these results, the EPR is computed, quantifying the fraction of the total electromagnetic energy stored in each nonlinear element. This quantity links the linear eigenmodes of the circuit to its nonlinear dynamics and provides a direct route to determine the contribution of each nonlinear element to the system Hamiltonian. The participation ratio, which serves as the bridge between linear and nonlinear circuit components, will be discussed in greater detail in the following sections.
@@ B.2 Full Hamiltonian of the System
where $m$ represents the electromagnetic mode, $\omega_{m}$ are the resonant frequencies of the modes and $\hat{a}_{m}^{{\dagger}}$ and $\hat{a}_{m}$ are the creation and annihilation operators for the given mode, respectively. If we are using linear elements in our circuit we can always describe the total energy with the Hamiltonian in Eqn. 13. Essentially, for every radiation field mode in the circuit we are associating with it a harmonic oscillator. Note that we have dropped the zero-point energy term, $\hbar\omega_{m}/2$, from Eqn 13 as is customary in renormalization to avoid infinite energies. To express the full Hamiltonian we combine the linear and nonlinear parts as follows:
@@
where the first term, introduced in Eqn. 13, includes the linearized Josephson junction. The second term represents the nonlinear contributions from the total number of Josephson junctions, $J$, present in the circuit. The nonlinear part of the full Hamiltonian can be expressed as:
$\mathcal{E}_{j}(\hat{\phi}_{j})^{\text{nlin}}$ $=-\frac{E_{J}}{24}\hat{\phi_{j}}^{4}+\frac{E_{J}}{720}\hat{\phi}_{j}^{6}$-... (15)
@@
where the angular brackets indicate to take the expectation value of the expression. Note that due to the energy balance between the inductive energy and capacitive energy stored in a given mode, the total inductive energy of the mode can be expressed as half the total energy. Then performing the calculation, where we take the expectation value with reference to a given Fock state $|n\rangle$, we end up with the result:
@@
where $S_{mj}\in\{-1,1\}$ is the sign of the junction and corresponds to the direction of current flow across the junction. From Eqn. 21 it can be seen that the participation ratio links the quantum properties of the circuit to its classical electromagnetic behaviour. Furthermore, all the components of the Hamiltonian are now fully specified and the system can now be solved by numerical diagonalization. The explanation provided in this section is a brief summary of the work in Ref. [minev_energy-participation_2021], for further details please refer to this paper.
@@ Appendix C Example: Transmon Coupled to a Resonator
The energy participation ratios for each mode are automatically calculated in post-processing by Palace. Recall that the EPR for a given mode is defined as the inductive energy stored in the junction divided by the total inductive energy in the mode as given in Eqn. (18). Additionally, the energy balance for a resonantly excited mode, $m$, due to time-averaging is equally split between inductive energy and electric energy [minev_energy-participation_2021]:
@@
Appendix D Computers Used in the Simulations
All Ansys and COMSOL simulations were run on a local HP DL380 computer which has a Dual Xeon E5-2697v4 CPU configuration with 768 GB of RAM. The Palace simulations were run on both this PC and on the Bunya cluster at UQ. Bunya possesses 113 AMD CPU nodes with 96 physical cores per compute node [bunya]. The results presented in this paper used 3 nodes each with 35 CPUs (105 CPUs total) where we balanced the trade-off between execution time and the queuing allocation time. To benchmark the speed increase, we increased the number of nodes from 1 to 10 (thereby, utilizing from 20 to 200 CPUs) to observe the resulting times for an eigenmode simulation. The simulation times shown in Fig. 6 were for a triple hanging resonator design with 24.5 million DoF. We fitted Amdahl’s law [Amdahl1967] to get an approximate single worker execution time of 88900 s and a single worker execution efficiency of 99.2%.
Figure 6: Simulation times when running Palace on a cluster. We benchmark the simulations times for an eigenmode simulation on a simple 3-resonator design with 24.5 million DoF. We fitted Amdahl’s law and found a fractional worker execution efficiency of 99.2% and a single worker execution time of approximately 88900 s.